If four times a number, increased by 2 is the same as three times a number increased by 9 what is the number?
Answer by Edwin McCravy(20056)

four times a number increased by 2 is the same as three times a number increased by 9
Replace the words "a number" by x
four times x increased by 2 is the same as three times x increased by 9
Replace the word "Four" by "4" and the word "three" by "3"
4 times x increased by 2 is the same as 3 times x increased by 9
Replace the words "4 times x" by "4x" and the words "3 times x" by "3x"
4x increased by 2 is the same as 3x increased by 9
Replace the words "increased by" by "+"
4x + 2 is the same as 3x + 9
Replace the words "is the same as" by "=":

4x + 2 = 3x + 9
4x+ 2 - 2 = 3x + 9 - 2
4x = 3x + 7
4x - 3x = 3x - 3x + 7
x = 7
The number is 7.
Edwin